OpenStudy (anonymous):

Factor this polynomial expression. x3 - 125

OpenStudy (ajprincess):

a^3-b^3=(a-b)(a^2+ab+b^2)

Parth (parthkohli):

Hint 1: 125 = 5^3 Hint 2: Use the formula (a + b)(a^2 + ab + b^2) to express the difference of cubes.

OpenStudy (ajprincess):

x^3-125=x^3-(5)^3

Parth (parthkohli):

(a - b)* I meant

OpenStudy (anonymous):

i cant do it bc i dont have a calculator

Parth (parthkohli):

You don't need a calculator lol

OpenStudy (anonymous):

i stil dont understand

Parth (parthkohli):

Never mind :/

OpenStudy (anonymous):

B. (x - 5)(x2 + 5x + 25)

OpenStudy (ajprincess):

in the formula given to u substitute x instead of a and 5 instead of b.

OpenStudy (ajprincess):

exactly right.

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

yay:)

OpenStudy (ajprincess):

:)